OpenStudy (anonymous):

2^{4/l \log_{2}x}=1/256

OpenStudy (anonymous):

2^{4/l \log_{2}x}=1/256

OpenStudy (anonymous):

\[2^{4\log_{2}(x)}=\frac{1}{256}\]

OpenStudy (anonymous):

yes, how do i solve it?

OpenStudy (anonymous):

please do help

OpenStudy (anonymous):

i didn't, i just tried to interpret what you wrote

OpenStudy (anonymous):

yes, so can you help me solve it?

OpenStudy (anonymous):

we have lots of choices here

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

so what are they

OpenStudy (anonymous):

\[256=2^8\] so \[\frac{1}{256}=2^{-8}\] so you could start with \[2^{4\log_{2}(x)}=2^{-8}\] then \[4\log_2(x)=-8\]then \[\log_2(x)=-2\] and so \[x=2^{-2}=\frac{1}{2^2}=\frac{1}{4}\]

OpenStudy (anonymous):

so it's .2500?

OpenStudy (anonymous):

or we could say \[2^{4\log_{2}(x)}=\frac{1}{256}\] \[2^{\log_{2}(x^4)}=\frac{1}{256}\] \[x^4=\frac{1}{256}\] \[x=\frac{1}{\sqrt[4]{256}}=\frac{1}{4}\] so

OpenStudy (anonymous):

they computer wouldnt accept the answer :(